OpenStudy (anonymous):

Solve . sin^2(2x)+sin^2(3x)+sin^2(4x)+sin^2(5x)=2

OpenStudy (anonymous):

\[2(\sin^2(2x)+\sin^2(3x)+\sin^2(4x)+\sin^2(5x)=2)\] \[1-\cos4x + 1-\cos6x+1-\cos8x+1-\cos10x=4\] \[\cos4x + \cos 6x + \cos8x+\cos10x=0\] summation gives \[\frac{\sin (\frac{4(2x)}{2})\cos(4x+\frac{3(2x)}{2})}{\sin 2x}=0\]

OpenStudy (anonymous):

And after summation?

OpenStudy (anonymous):

Solve for x

OpenStudy (anonymous):

Could you do the cos part ?

OpenStudy (anonymous):

cos(7x)=0 \[7x=2n \pi \pm \pi/2\]

OpenStudy (anonymous):

Thank you!

OpenStudy (jamesj):

Nicely done NSB.